### v1

# LiveResearchBench: A Live Benchmark for User-Centric Deep Research in the Wild

Abstract
Deep research—producing comprehensive, citation-grounded reports by searching and synthesizing information from hundreds of live web sources—marks an important frontier for agentic systems. To rigorously evaluate this ability, four principles are essential: tasks should be (1) user-centric, reflecting realistic information needs, (2) dynamic, requiring up-to-date information beyond parametric knowledge, (3) unambiguous, ensuring consistent interpretation across users, and (4) multi-faceted and search-intensive, requiring search over numerous web sources and in-depth analysis. Existing benchmarks fall short of these principles, often focusing on narrow domains or posing ambiguous questions that hinder fair comparison. Guided by these principles, we introduce LiveResearchBench, a benchmark of 100 expert-curated tasks spanning daily life, enterprise, and academia, each requiring extensive, dynamic, real-time web search and synthesis. Built with over 1,500 hours of human labor, LiveResearchBench provides a rigorous basis for systematic evaluation. To evaluate citation-grounded long-form reports, we introduce DeepEval, a comprehensive suite covering both content- and report-level quality, including coverage, presentation, citation accuracy and association, consistency and depth of analysis. DeepEval integrates four complementary evaluation protocols, each designed to ensure stable assessment and high agreement with human judgments. Using LiveResearchBench and DeepEval, we conduct a comprehensive evaluation of 17 frontier deep research systems, including single-agent web search, single-agent deep research, and multi-agent systems. Our analysis reveals current strengths, recurring failure modes, and key system components needed to advance reliable, insightful deep research.
1 Introduction
Deep research is the process of addressing complex, open-ended questions that require long-horizon, multi-step reasoning and planning. It involves exploring hundreds of live web sources and synthesizing them into comprehensive, citation-grounded, and structured outputs such as reports (OpenAI, 2025). This process closely mirrors how human researchers work—iteratively seeking, validating, and integrating evidence. Building systems capable of this marks an important frontier in AI: moving LLMs from simple chatbots to independent problem-solvers. However, progress is bottlenecked by a benchmarking and evaluation crisis: Which tasks are truly representative? How should ambiguities in query formulation be resolved? How can we reliably measure comprehensiveness (breadth and depth), consistency, factuality, and citation accuracy in long-form outputs while safeguarding against data contamination (Nguyen et al., 2025)?
Evaluating agentic tasks is challenging in general, and deep research is particularly difficult to benchmark due to its open-ended nature. To address this, we first propose four key principles for benchmarking deep research systems: tasks should be (a) user-centric, reflecting realistic information needs; (b) dynamic, requiring up-to-date information beyond parametric knowledge; (c) unambiguous, ensuring consistent interpretation across users; and (d) multi-faceted and search-intensive, requiring multi-hop search across diverse web sources and in-depth analysis. Recent benchmarks have introduced search-intensive tasks (Gou et al., 2025), but these are often static (and thus prone to contamination) (Bosse et al., 2025), domain-specific (Xu et al., 2025b; Patel et al., 2025), coarse-grained, or ambiguous (Du et al., 2025)—frequently omitting details such as the intended audience, required output format, or scope. Consequently, we still lack reliable tools to assess whether models are producing genuinely high-quality research or merely assembling plausible-sounding yet insufficient responses.
We address this gap with LiveResearchBench, a benchmark of 100 expert-curated queries paired with detailed checklists, constructed with over 1,500 hours of human labor. The tasks we benchmark span seven diverse domains (e.g., science, business, healthcare) and ten categories such as market analysis, literature review, policy evaluation, and topic exploration. Each query is multi-faceted: it specifies scope, audience, and format whenever appropriate, and many require real-time search at evaluation time. For example, one task asks for a comprehensive report on the evolution of artistic styles “up to the present {{date}},” requiring breadth across historical periods, depth of analysis on techniques and materials, and presentation as a formal academic report for undergraduate students (Figure 1). Compared to prior work, tasks in LiveResearchBench are realistic and unambiguous. It makes it possible to accurately gauge retrieval, synthesis, and reasoning in deep research.
While these queries capture more diverse scenarios, we must also ensure that the way we evaluate the resulting reports provides fine-grained insights into deep research agent capabilities. Evaluating the resulting long-form/open-ended reports poses unique challenges. For example, we cannot use simple string-matching approaches, while real-time queries imply that there is time-sensitive variation, without a fixed ground truths. Most critically, tasks are inherently multi-dimensional: we must assess coverage, reasoning, evidence use, and presentation quality. Human evaluation is costly and hard to scale, while naive LLM-as-a-judge setups produce inconsistent results.
To overcome these obstacles, we propose DeepEval, a comprehensive evaluation suite for research reports that spans both content-level and report-level metrics. Content-level metrics assess coverage and comprehensiveness, analysis depth, citation association, and factual accuracy. Report-level metrics evaluate presentation & organization, and logical & factual consistency. Each metric is paired with a tailored evaluation protocol (e.g., checklist-based, pointwise, pairwise, or rubric-tree), selected to ensure stable model assessment and high agreement with human evaluations. This design enables scalable, fine-grained, and reliable evaluation of system performance in dynamic research settings, particularly for open-ended, long-form reports.
We summarize the key contributions as follows:
We establish four core task design principles for deep research, grounded in user survey and comparative analysis, that ensure tasks are realistic, unambiguous, and evaluation-ready.
We introduce LiveResearchBench, a benchmark of 100 expert-curated, checklist-validated tasks across diverse domains and categories, explicitly designed for dynamic, multi-faceted research.
We propose DeepEval, a comprehensive evaluation suite for open-ended/long-form reports, covering six complementary dimensions and closely aligning with expert assessments.
We conduct a comprehensive evaluation of 17 state-of-the-art open-sourced and proprietary single- and multi-agent systems. Our analysis reveals systematic vulnerabilities and error patterns, showing that most systems act primarily as information collectors but struggle to produce insightful, well-reasoned reports.
2 Related Work
Single- and Multi-Agent Deep Research Systems. Equipping LLMs with tools—often referred to as agents—for grounded answers to complex queries has become a major research direction. Pioneering systems such as Deep Research using o3 (OpenAI, 2025) combined browsing and code execution and generate long reports. Many successors have since emerged (Xu & Peng, 2025; Java et al., 2025b). Architectures generally fall into two broad categories. Single-agent systems place all tool-use decisions in one model. Examples include function-calling LLMs (Yang et al., 2025; Agarwal et al., 2025) and ReAct-style agents (Yao et al., 2022). Multi-agent systems (MAS) instead coordinate specialized roles such as planner, researcher, and writer in a predefined workflow. Proprietary multi-agent systems such as Grok 4 Heavy (xAI, 2025a) and Manus (Manus, 2025) remain undisclosed, while examples of open-source systems include Open Deep Research (Alzubi et al., 2025), DeerFlow (ByteDance, 2025), and OpenManus (Liang et al., 2025). Despite rapid progress, it is still unclear which dimensions favor MAS vs. single-agent systems, what fundamental bottlenecks remain, and whether MAS truly deliver consistent gains (Kapoor et al., 2025). The lack of standardized benchmarks and evaluation frameworks further prevents rigorous comparison. Our work directly addresses this gap by developing a comprehensive evaluation framework, enabling rigorous comparison of single- and multi-agent systems for dynamic and multi-faceted research tasks.
Deep Research Benchmark. Several benchmarks have been proposed for evaluating deep research. DeepScholarBench (Patel et al., 2025) targets related-work generation; DeepResearch Bench (Du et al., 2025) include 100 short open-ended questions; Deep Research Bench (Bosse et al., 2025), LiveDRBench (Java et al., 2025a), and Mind2Web2 (Gou et al., 2025) focus on information-seeking tasks. While valuable, these benchmarks are typically domain-specific, ambiguous in query formulation, limited to short-form answers with low reasoning load, and static, making fair comparison difficult and preventing evaluation on evolving information. A comparison is shown in Figure 1. Recent efforts such as DeepResearchGym (Coelho et al., 2025) adopts short and static tasks, while ResearcherBench (Xu et al., 2025b) focuses narrowly on AI domain. In contrast, our LiveResearchBench provides user-centric, unambiguous, and time-varying tasks that are multi-faceted, search-intensive, and span diverse domains.
Evaluation of Long-form Answer. Recent work has moved beyond simple reference-based metrics. Some benchmarks emphasize coverage and factual grounding (Laban et al., 2023; 2024; Huang et al., 2024; Xu et al., 2025a), such as PROXYQA (Tan et al., 2024) (sub-question coverage) and LongFact (Wei et al., 2024) (claim-level verification). Others target structure and overall quality, including LongEval (Wu et al., 2025) (plan-based vs. direct generation) and HelloBench (Que et al., 2024) (human-aligned protocols). However, these efforts typically isolate a single dimension. In contrast, DeepEval offers a comprehensive suite: report-level metrics (presentation, consistency) and fine-grained dimensions (depth, checklist-based coverage, citation association, rubric-tree accuracy). Crucially, DeepEval is reliable and human-aligned: instead of holistic scores, we use structured checklists (e.g., factual errors as atomic tests) and aggregate results, yielding stable grades that closely track expert judgments.
3 LiveResearchBench
In this section, we present LiveResearchBench, our benchmark for evaluating agentic systems on challenging deep research tasks. We begin by outlining the key principles that govern task design (Section 3.1), then provide a high-level overview of the benchmark itself (Section 3.2). We then describe the data generation pipeline (Section 3.3) and conclude with our verification procedure to ensure benchmark quality (Section 3.4).
3.1 Task Design Principles
To inform the design of our deep research benchmark, we conducted a user survey with participants from diverse backgrounds and occupations, including enterprise professionals, academic researchers and students, and users performing everyday tasks (see Appendix G). The survey aimed to identify realistic and representative use cases of deep research systems across varied contexts. From this study, we distilled four key principles to guide task design:
User-centric: Tasks should reflect the needs of the intended audience. For example, a query posed by researchers may require technical jargon and fine-grained detail, while such complexity could overwhelm non-expert users.
Unambiguous: Tasks should be clearly specified, since the quality of the final output depends heavily on how well user requirements are interpreted.
Time-varying: Task requiring real-time search represent a large share of user queries. Such tasks are inherently resistant to data contamination from pretraining corpora, unlike static or time-bounded queries that risk becoming outdated or leaked as LLMs evolve.
Multi-faceted and search-intensive: Tasks should be sufficiently challenging, demanding multi-hop search across diverse web sources and in-depth analysis. This ensures evaluation of deep research systems extends beyond simple fact retrieval.
3.2 Benchmark Overview
LiveResearchBench is designed as a user-centric, unambiguous benchmark for deep research tasks that require broad, real-time information gathering. It consists of 100 expert-curated questions, each paired with detailed checklists, created through a six-stage curation pipeline (Figure 3) and validated via a five-step quality control process (Figure 4). It spans seven domains—Science & Technology, Economy & Business, Health & Wellbeing, Law & Governance, Society & Culture, Education & Knowledge, and Media & Entertainment—and ten task categories, including market analysis, technical support, decision support, policy and regulation, literature review, competitive analysis, pros-and-cons comparison, wide information search, and topic exploration (Figure 2). Example queries is shown in Figure 1, with additional examples in Appendix D111The full dataset will be open-sourced subject to institutional approval..
Compared to prior deep research benchmarks, LiveResearchBench tasks are designed to be more unambiguous and realistic (Figure 1). For example, Deep Research Bench (Bosse et al., 2025) includes search-intensive queries such as How many IM and GM account closures did chess.com report for 2024?, but these are static, time-bounded, and require minimal reasoning—limiting their ability to evaluate evolving real-world research. Likewise, DeepResearch Bench (Du et al., 2025) often omits crucial elements such as the target audience, content and format requirements, or a clearly defined scope. Such omissions leave queries open to multiple interpretations, even across runs of the same model, undermining evaluation reliability.
In contrast, tasks in LiveResearchBench are explicitly multi-faceted and grounded in user needs. For instance, the query in Figure 1 Create a comprehensive report about the evolution of artistic styles …Cover major historical periods from Ancient civilizations up to the present {{date}}. …Target the report for undergraduate-level students …Structure as a formal academic report …Include analysis of major artists, techniques, materials, and stylistic characteristics for each period and region. requires dynamic, up-to-date information search (with {{date}} replaced by the evaluation date), while also specifying the breadth and depth of coverage, the intended audience (undergraduate students), and the desired output format (a formal academic report). This level of clarity ensures consistent interpretation across models and imposes rigorous demands on information retrieval and in-depth reasoning.
3.3 Benchmark Construction
We curate the queries and checklists in LiveResearchBench through a six-stage pipeline (Figure 3). To capture realistic research needs, we start with user interviews and a crowd survey involving enterprise professionals, academics, and the general public. Participants were asked: What questions would you ask a deep research agent? (see Appendix G). The collected responses are then used to determine the task and domain distribution. Next, we hire domain experts from enterprise and academia to draft initial research questions aligned with these distributions.
To make the questions unambiguous and consistently interpretable across both models and human evaluators, we then prompt two top-tier deep research models, OpenAI o3 Deep Research and Gemini Deep Research, to generate possible clarification questions. Relying on a single model risks bias or omission, so using multiple models provides complementary perspectives. Human experts then review these clarifications, combine them with their domain expertise, and refine each query to establish a clear research scope and ensure a shared understanding of intent.
After the queries are finalized, we use GPT-5 to generate initial checklists. Each checklist decomposes a query into unit questions that test specific aspects of the required output. For example, for the query How large is the US enterprise AI services market in 2024–2025 (provide specific market size in USD and growth rates)?…, the corresponding checklist items include: 1) Does the report provide the market size for 2024 and 2025? 2) Does the report focus on the US market?… (9 checklist items in total). These checklists function as unit tests for the coverage dimension, verifying whether generated reports address all essential points in the query. This design enables consistent, reliable evaluation across different agentic systems.
3.4 Data Verification
A rigorous verification process by human experts is essential for benchmark construction. We adopt a five-stage verification pipeline (Figure 4). After collecting the research questions and corresponding checklists, we prepare detailed annotation guidelines for expert annotators (Appendix H). Following these guidelines, annotators independently assess each query and checklist item. Queries are labeled as either appropriate or not appropriate, while checklist items are labeled as appropriate, not appropriate, valid but not necessary, or needs modification. Next, a separate group of quality-control experts reviews these annotations and independently decides whether they agree or disagree with the initial assessments. A second quality-control stage then samples annotation results to detect potential systematic issues and to ensure that overall quality deviates by no more than 5% from the true data quality. Finally, a third group of experts performs the last round of verification and cross-checking. They resolve conflicts, refine the queries and checklists as needed, and finalize the dataset.
4 DeepEval: A Comprehensive Evaluation Suite for Deep Research
Evaluating long-form, open-ended generation poses unique challenges: responses are too diverse for string matching, real-time queries lack fixed ground truths, and tasks are inherently multi-faceted. To address this, we introduce DeepEval, a comprehensive evaluation suite for model-generated research reports. DeepEval assesses both report- and content-level quality across six complementary dimensions: ❶ Presentation & Organization, ❷ Factual & Logical Consistency, ❸ Coverage & Comprehensiveness, ❹ 0pt, ❺ Citation Association, and ❻ Citation Accuracy.
Evaluation Protocol.
As the metrics differ in granularity, no single evaluation protocol is sufficient. For metrics ❶-❺222❻Citation Accuracy requires verifying (statement, URL) pairs throughout the report. This task is naturally handled using the rubric tree protocol detailed in Section 4.2., we consider three options (Gu et al., 2025): (1) Checklist-based, where the judge evaluates the generated report against a checklist and provides a binary score (0/1) for each item in the checklist. (2) Pointwise (additive), where the judge aims to find concrete errors in the report according to the criteria. The final score is determined based on the number of errors. (3) Pairwise comparison, where the judge compares two reports side by side and selects the preferred one (or declares a tie). Based on preliminary experiments, we narrowed down the choice of protocol for each metric, and subsequently conducted a human alignment study on the selected protocols. As a result, we adopt Checklist-based for ❶ Presentation & Organization and ❸ Coverage & Comprehensiveness; Pointwise (additive) for ❷ Factual & Logical Consistency and ❺ Citation Association; Pairwise comparison for ❹ 0pt. These protocols achieve high agreement with human judgments (see Appendix B).
Remarks: Another possible evaluation protocol is to prompt an LLM judge to assign a single rating within a fixed range (e.g., 0-10), guided by rubrics that define the meaning of each score. However, our initial study showed that this approach is unsuitable for any of our metrics, yielding less than 60% agreement with human judgments even when using SoTA LLM judges (Gemini-2.5 Pro and GPT-5). Upon closer examination of the judge responses, we found that the judge models often produced conflated scores with shallow reasoning (e.g., “the report is overall well written because…”), rather than systematically identifying issues sentence by sentence. In other words, once a model commits to an overall positive assessment, it rarely examines the report in detail, leading to missed errors and inflated scores.
LLM-Ensemble-as-Judge. To mitigate inductive bias from relying on a single model, we conducted a pilot study to identify which top-tier models align best with human preferences. Among the candidates—Gemini 2.5 Pro, GPT-5, and Claude 4 Sonnet—results showed that Gemini 2.5 Pro aligned most closely, followed by GPT-5, while Claude 4 Sonnet exhibited inconsistent and low agreement with human experts (Appendix B). Based on these findings, we adopt a multi-judge ensemble protocol for all LLM-as-a-Judge evaluations. Specifically, we use Gemini 2.5 Pro and GPT-5 as independent judges and report final scores as the average of their assessments.
4.1 Coarse-grained: Report-level Metrics
❶ Presentation & Organization. Poorly organized reports or those with frequent grammatical errors undermine user trust. We evaluate presentation quality with a checklist of 10 common error patterns from our pilot study (Table 3), covering structure, grammar, citations, duplicates, and formatting. Each item is scored in binary form , and we report the average success rate across items and an ensemble of judge models as this configuration best aligns with human judgment—achieving agreement in 98.3% of cases (Appendix B). Examples of checklist items include: (1) Does the report contain any grammar or spelling errors? (2) If figures or tables are included, do they present complete data or valid visual elements? (3) Does every reference entry correspond to at least one in-text citation? Full checklists and prompts are provided in Appendix A.1.
❷ Factual & Logical Consistency. As reports grow longer, they are more likely to contain factual and logical inconsistencies, or even contradictory claims and numbers. This metric captures whether claims remain coherent and factually consistent throughout the document. We adopt the pointwise (additive) evaluation protocol based on the human alignment study, where the LLM judge is tasked with identifying as many substantive inconsistencies as possible. The final score is then computed mechanistically based on the number of issues detected, scaled from 0 to 100: a report with no contradictions receives a score of 100, while each additional inconsistency leads to a deduction. The ensemble of judges achieves a human agreement rate of 82%. Detailed prompts are provided in Appendix A.2.
4.2 Fine-grained: Content-level Metrics
❸ Coverage & Comprehensiveness. A strong research report should directly address all aspects of the multi-faceted query. To make evaluation consistent and interpretable, we leverage the human curated and verified checklists introduced in Section 3. Each checklist item functions as a unit test: the judge assigns a binary score (0/1) indicating whether the requirement is met, and we report the average success rate across all items and queries. The ensemble of judge models aligns with human judgment—100.0% of preferences favor either or both. The evaluation prompt is provided in Appendix A.3.
❹ Analysis Depth. Beyond coverage, a strong research report should provide substantive insights rather than superficial information gathering. We assess depth along five dimensions: (1) granularity of reasoning (causal chains), (2) multi-layer insights (trade-offs and implications), (3) critical evaluation (critiques and limitations), (4) analytical use of evidence, and (5) insight density. Ensemble judges conduct pairwise comparisons between the evaluated and baseline reports, assigning 1–5 scores to each along every dimension. To mitigate positional bias, we adopt the position-swap averaging method of Wang et al. (2023), scoring each pair twice with reversed order and averaging results. Dimension scores are summed, averaged across judges, and used to derive win rates: A win, B win, or Tie (if scores differ by ). Ties are excluded from the denominator. The human agreement rate is 92.5% with the judge ensemble. Detailed prompts can be seen in Appendix A.4.
❺ Citation Association. This metric checks whether all factual claims are linked to corresponding sources. For example, a statement like “medium-lift launch vehicles held 56.63% of the market in 2024” should include proper source URL. The LLM judge aims to flag claims not properly cited, and compute a score scaled 0–100: full coverage yields 100, with deductions for missing or misplaced citations, similar to the scoring mechanism in ❷. The human agreement rate is 85.9% with the judge ensemble. Detailed prompts can be seen in Appendix A.5.
❻ Citation Accuracy. Complementary to association, this metric evaluates whether citations genuinely support their claims. In citation-grounded reports, links may be hallucinated, irrelevant to the topic, or fail to substantiate the associated statement. To capture this, we employ an agentic judge with web access that performs a structured validation process: (1) identify each claim with its linked source URL, (2) check whether each URL is accessible, (3) if accessible, assess whether the content sufficiently supports the claim. This procedure is organized as a rubric tree in Figure 5.
As reports may contain extensive citations, verifying each link individually would be time-consuming and costly. To improve efficiency, we group all claims associated with the same link, requiring only a single web retrieval for the grouped claims. In addition, before checking whether the content supports a claim, we first perform a coarse-grained relevance check by examining the initial portion of the webpage (e.g., title or first few hundred tokens). For example, if the task concerns healthcare but the link directs to a technology article, we terminate early without further evaluation.
5 Main Results and analysis
Agentic Systems. We conduct a comprehensive evaluation of 17 state-of-the-art open-sourced and proprietary agentic systems, which can usually be grouped into three categories: (1) Single-agent systems with web search capabilities, including GPT-5 (OpenAI, 2025a) , GPT-4.1 (OpenAI, 2024), GPT-5-mini (OpenAI, 2025b), Gemini 2.5 Pro (DeepMind, 2025b), Gemini 2.5 Flash (DeepMind, 2025a), Claude 4 Sonnet (Anthropic, 2025a), Claude 4.1 Opus (Anthropic, 2025b), Perplexity Sonar Reasoning (Perplexity, 2025a), and Perplexity Sonar Reasoning Pro (Perplexity, 2025b); (2) Single-agent deep research systems, which feature extended reasoning depth and longer thinking time, including OpenAI o3 Deep Research (OpenAI, 2025c), OpenAI o4-mini Deep Research (OpenAI, 2025d), Perplexity Sonar Deep Research (AI, 2025b), Grok-4 Deep Research (Expert) (xAI, 2025b), and Gemini Deep Research (DeepMind, 2025c); (3) Multi-agent deep research systems, which coordinate a team of specialized agents to decompose complex queries. These systems typically employ a planner to dispatch sub-agents for browsing, searching, synthesis, and verification, while maintaining a large shared context across tasks. This category includes Manus (Manus, 2025), Grok-4 Heavy Deep Research (xAI, 2025a), Open Deep Research (AI, 2025a), and Deerflow+.
Deerflow+ is our enhanced implementation of Deerflow (ByteDance, 2025) that resolves frequent failure modes and early termination by adding inline citations, reference mapping, and robust long-context management. Compared to Deerflow, Deerflow+ eliminates token-limit failures, retains more retrieved information, and exhibits stronger evidence structuring and citation discipline. Implementation details are provided in Appendix C; Figure 13 shows side-by-side outputs for the same prompt in Figure 1.
Remarks: Non-reasoning models and multi-agent systems that do not excel at generating long-form reports are excluded due to poor performance in our pilot study. For each system, we configure settings to maximize performance. For example, setting output length to the maximum token limit (128,000 for GPT-5 models) and enabling retries for low-quality or failed generations. For open-source systems like Deerflow+ and Open Deep Research, we use GPT-5 as the backbone model.
[…]
Avg | ||||||||
[…]
| GPT-5 | 71.6 | 68.3 | 83.4 | 69.0 | 73.1 | ||||||||
| GPT-4.1 | 66.0 | 65.9 | 63.6 | 66.9 | 65.6 | ||||||||
| GPT-5-mini | 61.4 | 66.9 | 80.5 | 62.2 | 67.7 | ||||||||
| Gemini 2.5 Pro | 51.9 | 76.5 | 73.1 | 44.9 | 61.6 | ||||||||
| Claude 4 Sonnet | 81.9 | 67.3 | 49.2 | 50.8 | 62.3 | ||||||||
| Claude 4.1 Opus | 81.6 | 67.5 | 50.8 | 47.2 | 61.8 | ||||||||
| Perplexity Sonar Reasoning | 82.1 | 73.0 | 40.7 | 61.7 | 64.4 | ||||||||
| Perplexity Sonar Reasoning Pro | 79.6 | 71.9 | 46.7 | 65.0 | 65.8 | ||||||||
[…]
| OpenAI o3 Deep Research | 71.3 | 64.2 | 85.0 | 30.9 | 62.9 | ||||||||
| OpenAI o4-mini Deep Research | 74.3 | 62.3 | 78.6 | 32.1 | 61.8 | ||||||||
| Perplexity Sonar Deep Research | 83.5 | 67.4 | 65.5 | 52.1 | 67.1 | ||||||||
| Grok-4 Deep Research | 69.1 | 57.4 | 86.3 | 64.7 | 69.4 | ||||||||
| Gemini Deep Research | 62.1 | 63.0 | 75.8 | 64.6 | 66.4 | ||||||||
[…]
| Manus | 75.0 | 63.1 | 73.3 | 53.8 | 66.3 | ||||||||
| Grok-4 Heavy Deep Research | 75.9 | 59.4 | 89.3 | 64.7 | 72.3 | ||||||||
| Deerflow+ (w. GPT-5) | 78.8 | 69.9 | 61.6 | 81.4 | 72.9 | ||||||||
| Open Deep Research (w. GPT-5) | 81.0 | 71.3 | 65.3 | 77.2 | 73.7 | ||||||||
Obs.❶ Longer reports are common, but not consistently better. Figure 6 shows the distribution of report lengths across models and tasks. Single-agent deep research systems backed by strong base LLMs (Gemini Deep Research, OpenAI o3 Deep Research) produce substantially longer reports than both multi-agent systems and single-agent systems with web search. Yet, multi-agent pipelines do not automatically yield longer reports. For example, Deerflow+ and Open Deep Research rely on the final synthesizer to produce the report given searched findings, which constrains length. We also observe task-level variation on report length is smaller than model-level variation. Length can also be inflated by formatting choices rather than substance. Some systems emit very long, provider-specific redirect URLs in citations (e.g., “grounding” redirect URLs for Gemini 2.5 Pro), or duplicate links both inline and in the bibliography (OpenAI o3 Deep Research), effectively counting the same URL twice. Since we use shorten links and cite once for multi-agent systems for Deerflow+ and Open Deep Research, they yield fewer URL tokens. Hence, token counts reflect citation handling as much as content. As we show later, longer reports do not necessarily translate into higher quality; improvements in citation discipline and factual consistency are not monotonically related to length.
Obs.❷ Models struggle most with citation correctness and formatting, rather than surface fluency. Across systems, we repeatedly observe failure modes that are trivial for humans but challenging for current agents, such as mismatched in-text citations and references (Figure 14), missing or incomplete URLs (Figure 15), inconsistent citation formats (Figure 16), uncited references appearing in the bibliography (Figure 17), broken or incomplete table formatting (Figure 18), and out-of-order reference numbering (Figure 19). These recurring issues motivated the design of DeepEval’s Citation Association dimension (Section 4), and foreshadow the quantitative gaps we analyze next.
Table 1 summarizes performance across four DeepEval dimensions. Below we highlight key insights.
Obs.❸ Multi-agent families lead on average. At the system level, Open Deep Research achieves the highest average (73.6), followed by GPT-5 (72.7) and Deerflow+. Averaged by family, multi-agent deep research systems (69.5) outperform both single-agent web (62.8) and single-agent deep research, indicating that collaborative decomposition and aggregation provide reliable gains.
Obs.❹ Single web agent excels in consistency, multi-agent systems in association, while single-agent deep research lags. Single-agent web models are strongest on Factual & Logical Consistency (69.7 avg), benefiting from a single persistent memory stream that maintains reasoning continuity. Gemini 2.5 Pro achieves the highest consistency score overall (76.5), with Perplexity Sonar Reasoning and Perplexity Sonar Reasoning Pro close behind (73.0 and 71.9). For example, single web agents such as Gemini 2.5 Pro maintain coherence because the absence of inter-agent handoffs reduces the risk of logical drift. Similarly, Perplexity Sonar Reasoning benefits from producing relatively concise reports, which lowers the probability of factual or logical mistakes while keeping reasoning aligned. Multi-agent systems dominate Citation Association (61.9 avg), where explicit steps for aligning sub-agent outputs with citations or dedicated citation agents yield cleaner claim–evidence links. By contrast, single-agent deep research models perform worst on Citation Association, primarily because many key statements remain uncited and URLs are often mismatched. For example, OpenAI o3 Deep Research often produces long reports but leaves key statements/facts uncited, while Grok-4 Heavy Deep Research often cites URLs that fail to support the associated claims. Another common recurring issue is the use of fictional or non-existent links, as seen in Manus.
Obs.❺ Multi-agent systems lead presentation, but surface polish does not imply grounded quality. Multi-agent systems achieve the highest presentation scores (77.7 avg.), with Open Deep Research (81.0) and Deerflow+ (78.8) producing the most polished outputs. Perplexity and Claude models also perform strongly on this dimension, with scores above 79.5. However, presentation quality is weakly correlated with citation association or consistency: for example, Manus and Grok-4 Heavy Deep Research show good presentation (75.0, 75.9) but much lower citation association (53.8, 64.7), while Claude 4 Sonnet (81.9) and Perplexity Sonar Reasoning (82.1) similarly rank high in presentation yet fall to 50.8 and 61.7 in association, respectively. In contrast, Gemini 2.5 Pro performs poorly on presentation (51.9) but achieves the highest consistency score (76.5). In other words, visually well-organized reports may still contain inconsistent claims.
Obs.❻ Coverage benefits from specialization, but scaling retrieval scope and system complexity strain memory capacity. On average, single agents like GPT-5 series and the Grok family perform strongly on Coverage & Comprehensiveness, with Grok-4 Heavy Deep Research achieving the highest score (89.3). By contrast, Deerflow+ (61.6) and Open Deep Research (65.3) lag despite strong citation association. The bottleneck arises as the search context scales: even hundreds of retrieved webpages can quickly exceed the available context window. In multi-agent systems, increasing planning steps and subagents can easily push retrieval into thousands (3,000), where the problem compounds. Without robust mechanisms to label related content, track what has already been processed, and compress information without discarding key evidence, systems lose critical coverage. Thus, the open challenge is not the synthesizer alone, but the design of system-wide memory architectures and compression strategies that allow agents to retain breadth while maintaining fidelity.
Obs.❼ Most systems are deep searchers, not deep researchers. Figure 7 benchmarks analysis depth as win rate over Open Deep Research (with GPT-5 as backbone). Only Deerflow+ and Gemini Deep Research exceed ODR. Models that generate very long reports (e.g., OpenAI o3 Deep Research) often fail to synthesize across sources; Grok-4 Heavy Deep Research, though best on Coverage & Comprehensiveness, barely outperforms ODR in depth. This suggests that many systems behave more like information collectors and organizers than writers of evidence-grounded, argumentative reports—functioning as deep searchers rather than deep researchers that provide in-depth analysis. We provide further analysis with human judgment in Appendix B.
Obs.❽ Even SoTA systems are far from citation error-free. Using our rubric-tree framework, we evaluate the top performers from Table 1 (GPT-5, Grok-4 Deep Research, and Open Deep Research) on the two most search-intensive tasks: market analysis and wide information search. The results are shown in Table 2, where for each agentic system and task category, we report the average number of E1, E2, and E3 errors in the generated reports. We can see that all models produce non-trivial citation errors. In wide information search, most errors stem from unsupported claims (claims not verifiable from the cited link) rather than invalid or irrelevant URLs333E1 and E2 are counted at the URL level (each problematic URL contributes one error). E3 is counted at the statement level: all citations attached to a claim are expected to support it (e.g., if a claim about vaccination policy cites [5,10,12,24,25], the E3 error count is incremented by one for every source not supporting the claim., highlighting that hallucinations persist even with web access. The problem is worse for market analysis, where all systems generate large numbers of unsupported claims (Open Deep Research averages 91.9 errors per report), underscoring citation accuracy as a persistent bottleneck in deep research.
[…]
Total | ||||||
[…]
| GPT-5 | 4.2 | 1.7 | 13.3 | 19.2 | ||||||
| Grok-4 Deep Research | 6.8 | 6.8 | 33.4 | 47.0 | ||||||
| Open Deep Research | 5.0 | 5.2 | 19.7 | 29.9 | ||||||
[…]
| GPT-5 | 11.1 | 10.1 | 43.8 | 65.0 | ||||||
| Grok-4 Deep Research | 6.3 | 6.4 | 61.5 | 74.2 | ||||||
| Open Deep Research | 11.9 | 11.6 | 68.4 | 91.9 | ||||||
Obs.❾ Coherence, verifiability, breadth, depth trade off under context limits. Single web agent systems favor coherent reasoning; multi-agent systems achieve stronger association and presentation; and specialized designs lead in coverage. Yet no system achieves all four. Sustaining both high coverage and depth will require: (i) long-horizon memory with update operations, (ii) hierarchical information compression without information loss, and (iii) explicit synthesis/argumentation modules. Effective compression should address two challenges: (1) when information is partially redundant, the system should merge overlapping content without discarding unique evidence; and (2) when all information is related, compression should instead operate by assigning importance levels to retain critical details within the context limit. Because different users may emphasize different aspects of information, such representations must be both importance- and preference-aware, capturing what is essential for reasoning or verification while adapting to user priorities. Without these capabilities, scaling retrieval primarily increases report length rather than reliable, user-aligned insight.
6 Conclusion
We introduced LiveResearchBench, a benchmark of 100 expert-curated, dynamic tasks for deep research, and DeepEval, a comprehensive evaluation suite spanning six dimension for citation-grounded reports. Our evaluation of 17 leading systems shows that while agents can gather and organize information, they struggle with citation reliability and analytical depth. LiveResearchBench and DeepEval establish a rigorous foundation for benchmarking and point toward future advances in memory, compression, and synthesis as key to enabling truly insightful deep research.
Appendix
Appendix A Evaluation Details
A.1 Presentation & Organization
Grading prompt and checklist.
The prompt for presentation and organization is shown in Figure 8, and the corresponding checklist questions are provided in Table 3. The same checklist is used for all research questions.
| No. | Checklist Questions for Report Presentation Quality |
| 1 | Does the report present a clear, coherent, and logically ordered structure so that the organization is easy to follow and directly addresses the research question? |
| 2 | Does the report contain zero grammar and spelling errors? |
| 3 | Does every entry in the reference list correspond to at least one in-text citation? |
| 4 | Does every in-text citation have a corresponding entry in the reference list? |
| 5 | Is there exactly one “References” (or “Bibliography” / “Sources”) section, and are its entries sorted according to a single, consistent scheme? |
| 6 | Is a single, consistent citation style used throughout the entire document? |
| 7 | Are all in-text citations placed logically at the end of a clause or sentence, without interrupting grammatical flow? |
| 8 | If the report includes figures or tables, does each one contain complete data or a valid visual element? (If none are included, the report automatically passes this test.) |
| 9 | Is the formatting correct and consistent? For example: (a) If delivered in Markdown, are proper heading levels (#, ##, etc.) used instead of plain text for section titles; (b) if Markdown tables are included, is their syntax valid and renderable? |
| 10 | If the citations are numbered, are there no skipped numbers (e.g., [23], [25], [26] with [24] missing) and no duplicates (two different sources assigned the same number, or one source assigned multiple numbers)? |
A.2 Factual & Logical Consistency
Grading prompts and rubrics.
The grading prompt for factual and logical inconsistency is provided in Figure 9. The scoring rubrics are provided in Table 4. The scores for Factual & Logical Consistency are inversely correlated with the number of detected inconsistencies. Users can adjust the penalty weight assigned to each error depending on the evaluation need.
[…]
| 100 | 0 | Perfect – No contradictions or inconsistencies detected |
| 90 | 1–2 | Excellent – Very minor inconsistencies |
| 80 | 3–4 | Very Good – Few minor inconsistencies |
| 70 | 5–6 | Good – Some inconsistencies |
| 60 | 7–8 | Above Average – Several inconsistencies |
| 50 | 9–10 | Average – Multiple inconsistencies |
| 40 | 11–12 | Below Average – Many inconsistencies |
| 30 | 13–14 | Poor – Extensive inconsistencies |
| 20 | 15–17 | Very Poor – Pervasive inconsistencies |
| 10 | 18+ | Unacceptable – Overwhelming inconsistencies |
A.3 Coverage & Comprehensiveness
Grading prompt and checklist.
We present the prompt for evaluating Coverage & Comprehensiveness in Figure 10. An example coverage checklist is provided in Table 5. Note that the checklist for every research question is different (Sections 3.3 and 3.4).
[…]
| 1 | Does the report cover major historical periods from Ancient civilizations up to the present (evaluation date)? |
[…]
| 3 | Does the report provide analysis of key regional or national characteristics for European, East Asian (including China and Japan), Islamic, and Indigenous American art? |
[…]
| 5 | Does the report analyze how political, social, religious, and technological factors influenced artistic development? |
[…]
A.4 0pt
Grading prompt.
A.5 Citation Association
Grading prompts and rubrics.
We provide the grading prompt for citation association in Figure 12. The scoring rubrics are provided in Table 6. As the scores are determined by the number of claims not properly cited, users can adjust the penalty weight assigned to each error depending on the use case.
[…]
| 100 | 0 | Perfect – All major claims/facts have citations; fully traceable |
| 90 | 1–2 | Excellent – Very few claims lack citations; minimal impact |
| 80 | 3–4 | Good – Few claims lack citations; minor omissions |
| 70 | 5–6 | OK – Some claims lack citations |
| 60 | 7–8 | Above Average – Several claims lack citations |
| 50 | 9–10 | Average – Many claims lack citations; significant association issues |
| 40 | 11–12 | Below Average – Most claims lack citations |
| 30 | 13–14 | Poor – Extensive uncited claims; report poorly supported |
| 20 | 15–17 | Very Poor – Overwhelming lack of citations |
| 10 | 18+ | Unacceptable – Report largely untraceable |
Appendix B Alignment of LLM Judges with Human Experts
As deep research reports are often long-form and open-ended, we employ LLM judges to evaluate different aspects of the reports. To ensure the reliability, it is crucial that the judgments of LLMs align closely with those of humans. Recall that we use the following evaluation protocols: Checklist-based for ❶ Presentation & Organization and ❸ Coverage & Comprehensiveness; Pointwise (additive) for ❷ Factual & Logical Consistency and ❺ Citation Association; Pairwise comparison for ❹ 0pt. In the following, we discuss in more detail the procedure and results of our human alignment experiments on the chosen protocol for each metric.
Presentation & Organization.
We observe 67.8% agreement across the three LLM judges (Gemini 2.5 Pro, GPT-5, and Claude 4 Sonnet). In particular, Gemini 2.5 Pro and GPT-5 exhibit strong pairwise agreement (83.6%). To measure alignment with human judgment, we randomly sample 300 questions. For each question, a human expert is presented with the (Query, Report) pair and independently make binary decisions against the checklist without seeing LLM judge outputs. Subsequently, the human expert reviews the responses and rationales from the three LLM judges and asked to indicate their preference, choosing one judge, two jointly, declaring a tie, or none preferred. The human agreement rate for Gemini 2.5 Pro or GPT-5 is 98.3% of cases (Gemini 2.5 Pro alone 95.0%, GPT-5 alone 85.0%). Based on these results, we adopt an ensemble of Gemini 2.5 Pro and GPT-5 as our final judges for Presentation & Organization where each question is independently scored by both judge models and we report the average results.
Coverage & Comprehensiveness.
The overall agreement rate across the three LLM judges (Gemini 2.5 Pro, GPT-5, and Claude 4 Sonnet) reaches 81.5%, with particularly strong consistency between Gemini 2.5 Pro and GPT-5 (89.4%). To further assess alignment with human experts, we select 72 checklist questions and follow the same evaluation procedure used for Presentation & Organization. The human evaluator first provides independent judgments without access to model outputs, then reviews the LLM judges’ decisions and rationales, and finally indicates which model’s assessment best aligns with their own or none preferred. Across all 72 questions, the final human-aligned judgment corresponds to either Gemini 2.5 Pro or GPT-5 in 100.0% of cases, with no instance uniquely favoring Claude 4 Sonnet. Based on these results, we adopt an ensemble of Gemini 2.5 Pro and GPT-5 for Coverage & Comprehensiveness evaluation, where each question is independently scored by both judge models and the final result is averaged.
0pt.
78.8% of the judgments are consistent across the candidate judges (Gemini 2.5 Pro, GPT-5, and Claude 4 Sonnet). In particular, Gemini 2.5 Pro and GPT-5 agree in 89.6% of cases. To evaluate alignment with human judgment, we randomly select 40 samples (agreement set) where all judges agree and 40 samples (disagreement set) where they disagree. A human expert is presented with triplets (Query, Model A output, Model B output) and asked to make independent judgments (A wins, B wins, or a tie) without seeing the LLM judges’ decisions. For the disagreement set, the human expert also review the decision and the reasoning of each judge. Next, human experts indicate their preference among the judge candidates, choosing either a single judge, a joint preference for two judges (e.g., Gemini 2.5 Pro and GPT-5), a tie, or none preferred. On the agreement set, humans agree with the LLM judges 95% of the time. In the disagreement set, we have similar observations as in other metrics, where humans prefer Gemini 2.5 Pro or GPT-5 in 92.5% of cases.
Consistency and Citation Association.
Our preliminary analysis shows that the pointwise (additive) scheme for evaluating factual and logical inconsistency as well as citation association achieves substantially higher human agreement rates than direct scoring (with GPT-5 and Gemini-2.5 Pro ensemble as the judges). Specifically, for consistency, the agreement rate with human judgments is 82% over 128 inconsistencies identified by GPT-5 and Gemini-2.5 Pro. For citation association, the agreement rate reaches 85.9% over 128 issues identified by these judges. Note that our default weighting scheme decreases the score for every two issues detected. Therefore, the actual alignment of final scores with human judgment would be even higher.
Appendix C Deerflow+
To broaden our multi-agent comparison and study cross-system differences, we included Deerflow. However, in our setting, the vanilla system was brittle on long, search-heavy queries and frequently terminated with the error max_total_tokens (xxx) must be less than or equal to max_seq_len (xxx). Inspection indicated unchecked context growth across agents, retrieval traces, and retries. We also observed that the absence of inline-citation control made claims harder to anchor to sources and increased hallucination risk for the same set of retrieved pages. Therefore, adding inline citations both simplifies verification and reduces hallucinations, improving overall report quality.
Stabilization.
We therefore introduce Deerflow+, an enhanced variant that adds a lightweight context manager and inline-citation support. For stabilization, each agent is equipped with a context manager that (i) estimates tokens and enforces a total context budget before every model call, (ii) reserves response capacity via an input/output split, (iii) compresses histories and large research payloads with progressive, recency-preserving truncation, and (iv) supports per-agent budgets so retrieval-heavy roles compress more aggressively than synthesis roles. The mechanism is model-agnostic and transparent to the agent logic.
Citation Control.
We added inline-citation support (the original system only emitted a terminal bibliography) and a validation pass that enforces one-to-one mapping between in-text citations and reference entries while checking numbering continuity and duplicates.
Results.
With these changes, Deerflow+ completed the full evaluation suite without token-limit failures and produced higher-quality reports: better retention of retrieved evidence, improved formatting and factual consistency, and more reliable performance on presentation checks tied to citation management, particularly P4 (Citation Completeness) and P9 (Format Consistency) in Figure 22. Overall, these fixes substantially reduce early termination and bring Deerflow on par with other multi-agent systems in our experiments. Figure 13 shows example outputs for the same prompt in Figure 1; Deerflow+ exhibits higher information retention, stronger evidence structuring with inline citations, and no token-limit failures.
(a) Deerflow (vanilla)
(b) Deerflow+ (ours)
Figure 13: Qualitative outputs for the same prompt in Figure 1 from Deerflow (a) and Deerflow+ (b). Our Deerflow+ retains more retrieved information, presents stronger evidence structuring and citation discipline, while avoiding token-limit failures seen in the vanilla version.
Appendix D LiveResearchBench Demonstration
We provide concrete queries in LiveResearchBench across different task categories in Table 7, Table 8, and Table 9.
Category
Query
Competitive Analysis
Compare Spotify, Apple Music, Youtube Music, Amazon Music, Pandora, and Tidal on music discovery features (including algorithmic recommendations, curated playlists, and social discovery), audio quality (with reference to both technical specifications and perceived listening experience), and podcast selection (considering global availability), user interface design and ease of navigation, music library size and catalog completeness, exclusive content and early releases, pricing structures for individual and family plans, artist compensation and payout rates, and compatibility with high-end audio equipment as of now.
Literature Review
Help me produce a detailed academic report (literature review) on the evolution of evaluation practices for single-agent and multi-agent systems (based on large language models). Focus on the works published from 2023-2025 (including high-quality published papers, technical reports, and preprints). The report should include the following dimensions: (1) Benchmarks and sandbox environments: How do benchmarks for single-agent systems differ from those for multi-agent research in terms of task diversity, scalability, and realism? What are the trade-offs made to create the benchmark, their strengths and limitations? (2) Metrics: What families of evaluation metrics are most frequently applied in single-agent systems versus multi-agent systems, and what trade-offs do they entail? Does the metric rely on LLM-as-judge, agent-as-judge, human, or string matching? (3) Implementation and cost: How is the benchmark/sandbox implemented? Is human annotation involved in creating the benchmark or sandbox and how is human annotation involved (if applicable)? Is the benchmark/sandbox scalable and reliable? (4) Future Directions: What concrete research directions and design principles can advance the development of unified, generalizable evaluation pipelines?
Market Analysis
What has been the size, growth rate, and geographic distribution of the COVID-19 vaccine market between 2021 and 2025? What were the key demand phases (initial rollout, booster campaigns, variant-specific vaccines), and how have public health policies and supply constraints influenced vaccine adoption globally? How do Moderna, Pfizer-BioNTech, Johnson & Johnson, and Novavax compare in terms of global dose distribution volumes, clinical efficacy (trial versus real-world data), and manufacturing capacity? What flagship vaccines define their market strategies, and how is the market expected to evolve considering waning demand and shifting public health priorities?
Table 7: Example queries in LiveResearchBench (Part I).
Category
Query
Decision Support
In adults aged 30–70 with limited weekly training time ( min/week; with a min/week subgroup), what is the comparative effectiveness of high-intensity interval training (HIIT; including low-volume/sprint-interval variants), moderate-intensity continuous training (MICT), and hybrid HIIT+MICT protocols on (a) cardiorespiratory fitness (VO2max/VO2peak, mLkgmin-1), (b) cardiometabolic outcomes (HbA1c/fasting glucose, HOMA-IR, blood pressure, lipids, adiposity/liver fat), and (c) patient-centered outcomes (adherence, enjoyment, adverse events)? How are these protocols defined and evaluated across international randomized trials and meta-analyses (2018–2025), with supportive context from observational studies and major guidelines (e.g., ACSM/WHO), and which progression/maintenance plans (session frequency, intensity prescription, minimal effective dose; supervised vs. home-based delivery) sustain gains over 6–12 months? Consider moderators including sex, age band (30–49 vs. 50–70), baseline fitness, and cardiometabolic risk (overweight/obesity, prediabetes/metabolic syndrome, hypertension), as well as medication status.
Policy & Regulation
I work on information security and am responsible for incident disclosure, compare CISA’s CIRCIA regulations, the SEC’s cybersecurity disclosure rules, and sector-specific mandates under NERC CIP. Help me provide a cross-sector analysis (energy, finance, healthcare, and other critical infrastructure), focusing on the most current compliance challenges from the past 1-2 years. Analyze across reporting timelines, covered-entity definitions, enforcement mechanisms, and penalty structures. Provide: - A short description of each regulatory framework - A multi-column comparison table covering all dimensions - Multiple real-world examples of enterprise compliance challenges per framework - An analysis of strategic implications for U.S.-based and global enterprises operating under U.S. jurisdiction across multiple critical infrastructure sectors
Pros & Cons
Conduct a comprehensive analysis on whether social media is good for society, with a primary focus on Europe (including the UK, EU member states, and Nordic countries). The analysis should focus specifically on teenagers (ages 13-18) and their relationship with social media. For this report, define "good for society" across four dimensions: 1) mental health and well-being (e.g., anxiety, depression, self-esteem, and loneliness), 2) educational and developmental impacts (e.g., learning, attention span, and digital literacy), 3) social and civil outcomes (e.g., friendships, community engagement, political awareness, and misinformation exposure), and 4) safety and risk factors (e.g., cyberbullying, privacy, exploitation, and screen-time effects). In the report, I would like to have 6 pro arguments and 6 con arguments. Each pro or con argument should include: (1) A clear thesis statement (2) Multiple supporting points substantiated with credible evidence (including but not limited to statistics, direct quotes, case studies, expert testimony, industry or government reports, academic research, news sources, online forums, and social media), followed by in-depth reasoning and analysis. Prioritize data and research from the past 5 years where available (2020-2025).
Table 8: Example queries in LiveResearchBench (Part II).
Category
Query
Topic Exploration
As someone who created an online photo editing tool targeting social media marketers and small businesses, help me understand how AI is transforming the photo editing space and what I need to do to stay competitive. My tool currently offers basic editing features like cropping, filters, and text overlays, and I’m considering adding AI capabilities through third-party APIs rather than building in-house. I want to know (1) which specific AI features would be most valuable for social media marketers - that I could integrate without massive development costs, (2) whether AI will make traditional photo editing skills obsolete or if there’s still a place for manual editing tools alongside AI features, (3) how major players like Canva, Photoshop, and newer AI-first competitors are using AI to attract users, and (4) what the real costs and benefits are for integrating AI APIs into my platform, including monthly API fees, development timeline, server load, and impact on user engagement and premium conversions. Give me examples from photo editing tools serving similar markets, with specific AI features, costs, and results.
Wide Info Search
Identify U.S. and Canadian banks and broker-dealers that, according to publicly available sources, launched (i.e., production, not pilot) client-facing AI portfolio-management or financial-planning ‘copilots’ during 2023-2025, and compile a due-diligence table covering: institution; assistant/product name; launch month; customer tier; geography & languages; core AI stack; safety & compliance; integrations (e.g., custodian/portfolio accounting/CRM/planning tools, SSO, audit logging); access & pricing; KPIs (e.g., adoption, satisfaction, ROI) where available; plus links to the bank/OEM press release, product page or demo, and at least one independent wealth-tech or fintech-trade article confirming the launch.
Top Rankings
What are the top 20 gaming monitors released or available for purchase in 2025, based on refresh rate (Hz), response time (ms), and price per inch of screen size? Please assume all three criteria are weighted equally. No specific panel technologies or size requirements are set; consider all currently available monitors within this timeframe.
Table 9: Example queries in LiveResearchBench (Part III).
Appendix E Error Pattern Examples
In this section, we provide concrete examples of error patterns observed in our pilot study. This includes mismatched in-text citations and references (Figure 14), missing links or incomplete URLs (Figure 15), inconsistent citation formats (Figure 16), uncited references appearing in the bibliography (Figure 17), broken or incomplete table formatting (Figure 18), out-of-order reference numbering (Figure 19), embedded citations breaking text flow (Figure 20), and hallucinated information (Figure 21). These representative cases illustrate the diverse range of structural and citation-related issues that can undermine the clarity and overall reliability of generated research reports. These observations motivated the tailored evaluation metrics in DeepEval as described in Section 4.
Figure 14: Error pattern: mismatched in-text citations and references.
Figure 15: Error pattern: missing links (URLs).
Figure 16: Error pattern: inconsistent citation format.
Figure 17: Error pattern: references not mentioned in the report.
Figure 18: Error pattern: broken table format.
Figure 19: Error pattern: out-of-order references.
Figure 20: Error pattern: embedded citations breaking text flow.
Figure 21: Error pattern: hallucinated information.
Appendix F Information on the Annotation Team
For data annotation and verification, we worked with professional annotators from Salesforce Inc. All annotators hold advanced degrees in fields including English Literature, Social Science, Computer Science, and Finance, with 2–6 years of experience in data annotation and analysis. Their expertise spans multi-modal evaluation (text, image, and audio) as well as quality control for advanced annotation pipelines.
Appendix G User Survey
To inform the design of our deep research benchmark, we conducted an online user survey with participants
from diverse backgrounds and occupation. We group participants into three categories: enterprise professionals (including sales, marketing, legal, and software engineering; ), researchers (including professors, industrial researchers, postdocs, and Ph.D. students; ), and general users (the rest; ). Participants were asked: What questions would you ask a deep research agent?, and What aspects of the model’s response matter most to you?
We analyzed the collected responses and observed the following trends: (1) many questions are multi-faceted and require real-time search; (2) users value responses that capture both the surface-level request and the implicit intention of the query; and (3) they expect the final report to be tailored to the target audience, in content, format, and presentation style.
For the collected questions, we distilled them into ten representative task categories (Figure 2). Among these, the three most common were topic understanding (24.5%), wide information search (22.8%), and top rankings (16.7%), followed by market analysis (8.2%), technical support (6.5%), decision support (5.2%), policy & regulation (4.8%), literature review (4.3%), topic exploration (3.5%), pros & cons (2.0%), and competitive analysis (1.5%). In addition, we also analyzed the domain distributions based on the collected user responses: Science and Technology, Economy and Business, Health and Wellbeing, Law and Governance, Society and Culture, Education and Knowledge, Media and Entertainment. Together, these findings were instrumental in shaping the design principles and the curation of tasks in LiveResearchBench.
Appendix H Human Verification Guidelines
Goal.
Verify the quality of deep research queries and their corresponding checklists that aims to assess the comprehensiveness and coverage of the report.
Data Format
The CSV contains one row per grading criterion, with the following key columns:
[…]
Column
Description
qid
Query identifier.
Deep Research Query
The detailed deep research query.
Checklist
Individual grading questions (Multiple checklist items per qid).
Table 10: Key columns in the annotation CSV format.
You will be evaluating two components: (1) Deep Research Query Quality and (2) Checklist Quality.
H.1 Annotation Tasks
H.1.1 Task 1: Deep Research Query Quality
Rate each unique Deep Research Query according to Table 11:
[…]
Rating
Description
Appropriate
Query is well-structured, complete, clear, and answerable through research.
Appropriate with modification
Good foundation but needs improvement (please provide a modified version).
Not appropriate
Fundamentally flawed or not worth fixing.
Table 11: Rubric for rating deep research query quality.
Consider the following criteria:
Answerability: Can this question be reasonably answered using web search and analysis?
Clarity: Is it clear what information is being requested? Useful clarifiers include but not limited to temporal bounds, geographic scope, and specific entities.
Completeness: Is the query complete? Would you, as the user, add or remove specifics to make it more complete?
format: When appropriate, is the format and report organization requirement clear? This criterion may not apply to all questions—use it judiciously.
Target audience: When appropriate, is the query framed appropriately for its intended readers (e.g., technical experts, policymakers, or general public)? Does the wording match the expected background knowledge and level of detail of that audience? This criterion may not apply to all questions—use it judiciously.
H.1.2 Task 2: Grading Criteria Quality
Rate each Checklist item according to Table 12:
[…]
Rating
Description
Appropriate
Relevant, verifiable, and important.
Valid but not necessary
Correct but redundant or overly granular (please comment why; modify if needed).
Needs modification
Poorly phrased or too narrow (please provide a modified version).
Not appropriate
Irrelevant, unverifiable, or impossible (please comment why; modify if needed).
Table 12: Rubric for rating checklist quality.
Remark: If additional checklist items should be included, you should append the new checklist items for the corresponding qid.
When evaluating each checklist item, consider the following criteria:
Relevance: Does the criterion correspond to the requirements of the research question?
Verifiability: Can it be objectively verified from a research report?
Importance: Does it capture a meaningful aspect of the research question?
Non-redundancy: Does it avoid substantial overlap with other checklist items?
H.2 Concrete Examples
H.2.1 Task 1: Deep Research Query
Rating: Appropriate. The query is well-structured with clear sub-questions, specifies a concrete temporal scope (2024) and geographic focus (U.S.), mentions relevant entities (Tesla, BYD, etc.), and is answerable through standard market research. Note that the query is well scoped with with a business/market research context, and it doesn’t hinge on whether the target audience is, say, a policymaker, an investor, or the general public—all of those groups could use essentially the same data. Therefore, there is no need to further specify the target audience in this case.
Rating: Needs Modification or Inappropriate. The query is overly broad (“LLM technology” is unspecified), lacks temporal bounds, and sets vague expectations (“current trends” without context). In addition, it does not reflect the needs of a specific target audience—while researchers may require technical detail, the general phrasing here fails to indicate whether the query is meant for experts, practitioners, or the broader public.
H.2.2 Task 2: Checklist Items
Rating: Appropriate. This checklist item is directly relevant to the research question, as it captures the query’s explicit focus on the U.S. electric vehicle market. It is objectively verifiable, since one can check whether the report includes U.S.-specific market data. It addresses an important dimension of the query by ensuring geographic precision, and it does so without overlapping with other checklist items.
Rating: Valid but not necessary. This checklist item is correct and verifiable, since Tesla is explicitly part of the query. However, it is overly granular: focusing only on Tesla misses the broader comparison with other key entities such as BYD, Volkswagen, and Ford.
Comment: Redundant if there is already a broader criterion covering multiple entities.
Rating: Not appropriate. This checklist item is unverifiable in practice, as there is no authoritative ground truth for identifying all electric vehicle companies. It also sets an impossible standard, since completeness across the entire market cannot reasonably be expected from a single report. A more suitable criterion would focus on specific entities mentioned in the query (e.g., Tesla, BYD, Volkswagen, Ford) rather than requiring exhaustive coverage.
H.3 Remarks and Tips
Each deep research query should be rated once for its corresponding qid, while each checklist item should be rated individually. Several tips for annotation: (1) avoid checklist items that demand unverifiable completeness, such as those asking for “all” or “every” entity or fact; (2) watch for redundancy, where multiple criteria check essentially the same requirement; (3) be cautious of scope creep, in which a checklist item evaluates something not actually specified in the query; (4) exclude subjective judgments; checklist items should focus on verifiable facts rather than opinions.
Appendix I Results Breakdown
Figure 22: Fine-grained pass rates (P1–P10) for the Presentation & Organization metric. Most systems are reliable on P1 (Structure & Coherence) and near-saturated on P7 (Citation Placement) and P8 (Figures & Tables Validity). The largest dispersion appears in grammar, citation style/formatting, and reference alignment (P2–P6, P9, P10). Models from the same provider exhibit correlated patterns; pipeline design (e.g., multi-agent wrappers) can help citation-focused tests.
Figure 22 reports the breakdown of the ten presentation checklist items (Table 3) for each system. High-level organization is largely under control: models perform well on P1 (section-level structure), and straightforward behaviors such as placing citations at clause/sentence boundaries (P7) and ensuring figures/tables contain valid entries (P8) are near-saturated. The harder cases are the mechanics of style and bibliographic control: style/formatting (P2, P5–P6, P9) and citation management (P3–P4, P10) show wide variance across systems. We also observe provider-level regularities: sibling models tend to share strengths and weaknesses (e.g., on P6, GPT-5 / GPT-4.1 / GPT-5 Mini score 31%, 32%, 24%), likely due to shared data and formatting designs.
System design can shift these profiles: multi-agent wrappers around a common backbone often improve P4 by enforcing explicit aggregation/mapping from in-text citations to references, whereas numbering accuracy in P10 varies by pipeline, with some single-agent web systems exhibiting steadier numbering.
Overall, a consistent pattern emerges across families: high-level coherence and simple placement/validity are easy; citation correctness, style uniformity, grammar, and global formatting discipline remain the hard parts.